### v1

@@ Appendix C Proof of Theorem 1
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})$
$\sigma^{2}_{*}$ =$=\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c<t<n.\end{cases}$ (43)
$\rho_{*}$ =$=\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}+\frac{\sqrt{\left((n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\right)^{2}+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (44)
@@
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\geq|\mathcal{F}|\leq n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}_{*}+\rho_{*}(t-1)}\right)^{-1}$ (45)
Proof:  The proof consists of two main steps, namely, 1) optimizing the decoder at the server for any fixed joint distribution of $(\bm{Z}_{1},\dotsc,\bm{Z}_{n})$, i.e., any fixed $\sigma^{2}$ and $r$ (equivalently $\rho$), 2) determining the feasible regions of $\sigma^{2}$ and $r$ that satisfy the privacy constraint, 3) characterizing the overall minimum MSE by optimizing $\sigma^{2}$ and $r$. Step 1 is proved in Proposition 1. For step 2, note that the information available at the server increases with the number of responding users and the number of colluding users. Therefore, we consider the case where all $n$ users respond, and up to any $c$ users can collude with the server, to analyze the privacy constraint. Lemma  1 characterizes step 2. We analyze step 3 in two cases. In case 1, we assume no dropouts, i.e., $t=n$. In case 2, we assume that up to any $t<n$ users can dropout. The two cases are analyzed in Lemmas 2 and 3, respectively.
Let $D_{Z}^{G}(\sigma^{2},r)$ denote the multivariate Gaussian distribution (of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$) with following properties. $\bm{Z}_{i}\sim\mathcal{N}(\bm{0}_{d},\sigma^{2}\mathsf{I}_{d})$, $i\in[1:n]$, with the all zeros vector of size $d\times 1$ denoted by $\bm{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\bm{Z}_{i}$ is denoted by $Z_{i,k}$ for $i\in[1:n]$ and $k\in[1:d]$. $Z_{i,k}$’s are allowed to be correlated as $\mathbb{E}[Z_{i,k}Z_{j,k}]=r$ for $\forall i\neq j$, $k\in[1:d]$. That is, $[Z_{1,k},\dotsc,Z_{n,k}]^{T}\sim\mathcal{N}(\bm{0}_{n},\Sigma)$ for $k\in[1:d]$ with $\Sigma_{i,i}=\sigma^{2}$ for $i\in[1:n]$ and $\Sigma_{i,j}=r$ for $i,j\in[1:n]$, $i\neq j$. Moreover, let $\mathbb{E}[Z_{i,k}Z_{j,k^{\prime}}]=0$, $\forall i,j$, $\forall k\neq k^{\prime}$.
@@
The Gaussian mechanism $\mathcal{D}_{Z}^{G}$ with given $\sigma^{2}$ and $r$ satisfies the privacy constraint in Definition 1 for any $\mathcal{F}$ and $\mathcal{C}$ so long as:
$r^{2}(n-1)(c-1)+r(\sigma^{2}(n+c-2)-\sigma^{2}_{\epsilon,\delta}(n-2))+\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\geq 0,$ (46)
@@
$r\geq\begin{cases}\frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (47)
@@
Proof:  [Proof of Lemma 1] Recall that the privacy mechanism of each user is given by $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$, $i\in[1:n]$. Note that for user $i$, the added noise can be written in the following form, based on our noise generation protocol in Section 4.
@@
where $\mathbf{S}_{i,j}\sim\mathcal{N}(\mathbf{0}_{d},-r\mathsf{I}_{d})$ and $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},(\sigma^{2}+r(n-1))\mathsf{I}_{d})$. To analyze the privacy constraint in Definition 1, the information available to the server from all users except for user $i$, $i\in[1:n]\setminus\mathcal{C}$ when any $\mathcal{C}\in[1:n]$, $|\mathcal{C}|\leq c$ users are allowed to collude with the server is given by,
$\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in[1:n],j\neq i},\{\mathbf{x}_{k},\mathbf{N}_{k},\{\mathbf{S}_{k,j}\}_{\forall j}\}_{k\in\mathcal{C}}\}.$ (52)
@@
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C},$ (53)
$\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$, for any fixed $\mathbf{x}_{j}$, $j\neq i$ and $\forall\mathcal{A}$, simplifies to,
$\mathbb{P}(\hat{M}(\mathbf{x}_{i})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[1:n],j\notin\mathcal{C},j\neq i)$
$\qquad\leq e^{\epsilon}\mathbb{P}(\hat{M}(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[1:n],j\notin\mathcal{C},j\neq i)+\delta,\quad\forall i\in[1:n]\setminus\mathcal{C},$ (54)
where $\hat{M}(\mathbf{x}_{k})=\mathbf{x}_{k}+\sum_{j<k,j\notin\mathcal{C}}\mathbf{S}_{j,k}-\sum_{j>k,j\notin\mathcal{C}}\mathbf{S}_{k,j}+\mathbf{N}_{k}$. Let $\mathbf{Y}_{k}=\hat{M}(\mathbf{x}_{k})$, $\forall k\in[1:n]$. We first derive the conditional distribution of $\mathbf{Y}_{i}$ given $\mathbf{Y}_{j}$, $\forall j\notin\mathcal{C}\cup\{i\}$. Without loss of generality assume that $i=1$ and $\mathcal{C}=\{n-c+1,\dotsc,n\}$.
@@
The conditional distribution of $\mathbf{Y}_{1}$, given $\mathbf{Y}_{j}$, $j\in[2:n-c]$ for any fixed $\mathbf{x}_{j}$, $j\in[2:n-c]$ is given by,
@@
$\tilde{\mathbf{\mu}}$ $=\mathbf{x}_{1}+\left(r\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\otimes\mathsf{I}_{d}\right)\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (56)
$\tilde{\Sigma}$ $=\left(\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\right)\mathsf{I}_{d}$ (57)
@@
Proof:  [Proof of Claim 1] $\mathbf{Y}_{\ell}=\hat{M}(\mathbf{x}_{\ell})=\mathbf{x}_{\ell}+\sum_{j<\ell,j\notin\mathcal{C}}\mathbf{S}_{j,\ell}-\sum_{j>\ell,j\notin\mathcal{C}}\mathbf{S}_{\ell,j}+\mathbf{N}_{\ell}$, for all $\ell\notin\mathcal{C}$. Therefore, $\mathbf{Y}_{\ell}\sim\mathcal{N}(\mathbf{x}_{\ell},(\sigma^{2}+rc)\mathsf{I}_{d})$. The distribution of the $k$th component of $\mathbf{Y}_{\ell}$, across all $\ell\in\mathcal{C}$ is given by,
$\begin{pmatrix}Y_{1,k}\\ \vdots\\ Y_{n-c,k}\end{pmatrix}\sim\mathcal{N}\begin{pmatrix}\begin{pmatrix}x_{1,k}\\ \vdots\\ x_{n-c,k}\end{pmatrix}_{(n-c)\times 1},\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c)\times(n-c)}\end{pmatrix}$ (59)
for $k\in[1:d]$, as $\text{cov}(Y_{i,k},Y_{j,k})=-\mathbb{E}[S_{i,j}^{2}]=r$ for $i\neq j$ and $\text{var}(Y_{i,k})=-r(n-c-1)+\sigma^{2}+r(n-1)=\sigma^{2}+rc$, $\forall i$. From the above distributions and from the fact that $\text{cov}(Y_{i,k},Y_{j,k})=0$, $\forall i,j\in[1:n]\setminus\mathcal{C}$ and $\forall k\neq k^{\prime}$, we derive,
$\begin{pmatrix}\mathbf{Y}_{1}\\ \vdots\\ \mathbf{Y}_{n-c}\end{pmatrix}\!\sim\!\mathcal{N}\begin{pmatrix}\!\!\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{n-c}\end{pmatrix}_{\!\!\!(n\!-\!c)\times 1}\!\!\!,\begin{pmatrix}(\sigma^{2}+rc)\mathsf{I}_{d}&r\mathsf{I}_{d}&\dotsc&r\mathsf{I}_{d}\\ r\mathsf{I}_{d}&(\sigma^{2}+rc)\mathsf{I}_{d}&\dotsc&r\mathsf{I}_{d}\\ \vdots&\vdots&\vdots&\vdots&\\ r\mathsf{I}_{d}&r\mathsf{I}_{d}&\dotsc&(\sigma^{2}+rc)\mathsf{I}_{d}\end{pmatrix}_{\!\!\!(n\!-\!c)d\times(n\!-\!c)d}\end{pmatrix}$ (60)
@@
Based on this result, we have, $\mathbf{Y}_{1}|\mathbf{Y}_{j};j\in[2:n-c]\sim\mathcal{N}(\tilde{\mathbf{\mu}},\tilde{\Sigma})$ where,
$\tilde{\mathbf{\mu}}$ $=\mathbf{x}_{1}+(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})\left(\begin{pmatrix}(\sigma^{2}+rc)&r&\dotsc&r\\ r&(\sigma^{2}+rc)&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&(\sigma^{2}+rc)\end{pmatrix}_{(n-c-1)\times(n-c-1)}\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)^{-1}$
$=\mathbf{x}_{1}+\left(r\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}(\sigma^{2}+rc)&r&\dotsc&r\\ r&(\sigma^{2}+rc)&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&(\sigma^{2}+rc)\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\otimes\mathsf{I}_{d}\right)$
@@
$\tilde{\Sigma}$ $=(\sigma^{2}+rc)\mathsf{I}_{d}-(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})\left(\!\!\!\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}\!\!\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)^{-1}$

### v2

@@ Appendix C Proof of Theorem 1
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$\sigma^{2}_{*}$ =$=\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}\\ +\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c<t<n.\end{cases}$ (65)
$\rho_{*}$ =$=\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}\\ +\frac{\sqrt{\left((n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\right)^{2}+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (66)
@@
$\min_{\sigma^{2},\rho}$ $\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$=\left(1+\frac{t/d}{\sigma^{2}_{*}+\rho_{*}(t-1)}\right)^{-1}$ (67)
Proof:  The proof consists of three main steps, namely, 1) optimizing the decoder at the server for any fixed joint distribution of $(\boldsymbol{Z}_{1},\dotsc,\boldsymbol{Z}_{n})$, i.e., any fixed $\sigma^{2}$ and $\rho$, 2) determining the feasible regions of $\sigma^{2}$ and $\rho$ that satisfy the privacy constraint, 3) characterizing the overall minimum MSE by optimizing $\sigma^{2}$ and $\rho$. Step 1 is proved in Proposition 1. For step 2, note that the information available at the server increases with the number of responding users and the number of colluding users. Therefore, we consider the case where all $n$ users respond, and up to any $c$ users can collude with the server, to analyze the privacy constraint. Lemma  1 characterizes step 2. We analyze step 3 in two cases. In case 1, we assume no dropouts, i.e., $t=n$. In case 2, we assume that up to any $t<n$ users can dropout. The two cases are analyzed in Lemmas 2 and 3, respectively.
Let $D_{Z}^{G}(\sigma^{2},\rho)$ denote the multivariate Gaussian distribution of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ with following properties. $\boldsymbol{Z}_{i}\sim\mathcal{N}(\boldsymbol{0}_{d},\sigma^{2}\mathsf{I}_{d})$, $i\in\mathcal{U}_{all}$, with the all zeros vector of size $d\times 1$ denoted by $\boldsymbol{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\boldsymbol{Z}_{i}$ is denoted by $Z_{i,k}$ for $k\in[1:d]$. $Z_{i,k}$’s are allowed to be correlated as $\mathbb{E}[Z_{i,k}Z_{j,k}]=\rho\sigma^{2}=r$ for $\forall i\neq j$, $k\in[1:d]$. That is, $[Z_{1,k},\dotsc,Z_{n,k}]^{T}\sim\mathcal{N}(\boldsymbol{0}_{n},\Sigma)$ for $k\in[1:d]$ with $\Sigma_{i,i}=\sigma^{2}$ for $i\in[1:n]$ and $\Sigma_{i,j}=r$ for $i,j\in[1:n]$, $i\neq j$. Moreover, let $\mathbb{E}[Z_{i,k}Z_{j,k^{\prime}}]=0$, $\forall i,j$, $\forall k\neq k^{\prime}$.
@@
The Gaussian mechanism $\mathcal{D}_{Z}^{G}$ with given $\sigma^{2}$ and $r$ satisfies the privacy constraint in Definition 1 for any $\mathcal{U}$ and $\mathcal{U}_{col}$ satisfying $|\mathcal{U}_{col}|\leq c<t\leq|\mathcal{U}|$, so long as:
$r^{2}(n-1)(c-1)+$ $r(\sigma^{2}(n+c-2)-\sigma^{2}_{\epsilon,\delta}(n-2))$
$+\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\geq 0,$ (68)
@@
$r\geq\begin{cases}\frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}\\ +\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (69)
@@
Proof:  [Proof of Lemma 1] Recall that the privacy mechanism of each user is given by $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$, $i\in[1:n]$. Note that for user $i$, the added noise can be written in the following form, based on our noise generation protocol in Section IV.
@@
where $\mathbf{S}_{i,j}\sim\mathcal{N}(\mathbf{0}_{d},-r\mathsf{I}_{d})$ and $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},(\sigma^{2}+r(n-1))\mathsf{I}_{d})$. To analyze the privacy constraint in Definition 1, the set of random variables observed by the server from all users except for user $i$, $i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$ when any $\mathcal{U}_{col}\subset\mathcal{U}_{all}$, $|\mathcal{U}_{col}|\leq c$ users are allowed to collude with the server is given by,
$\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in\mathcal{U}_{all}\setminus\{i\}},\{\mathbf{x}_{k},\mathbf{N}_{k},\{\mathbf{S}_{k,j}\}_{\forall j}\}_{k\in\mathcal{U}_{col}}\}.$ (74)
@@
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,$ (75)
$\forall\mathcal{D}_{i},\mathcal{D}_{i}^{\prime}$, $\forall i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$, and $\forall\mathcal{A}$, simplifies to,
$\mathbb{P}(\hat{M}(\mathbf{x}_{i})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in\mathcal{U}_{all}\setminus\{\mathcal{U}_{col}\cup i\})$
$\leq e^{\epsilon}\mathbb{P}(\hat{M}(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in\mathcal{U}_{all}\setminus\{\mathcal{U}_{col}\cup i\})+\delta,$ (76)
$\forall\mathcal{D}_{i},\mathcal{D}_{i}^{\prime}$, $\forall i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$, and $\forall\mathcal{A}$ where $\hat{M}(\mathbf{x}_{i})=\mathbf{x}_{i}+\sum_{j<i,j\notin\mathcal{U}_{col}}\mathbf{S}_{j,i}-\sum_{j>i,j\notin\mathcal{U}_{col}}\mathbf{S}_{i,j}+\mathbf{N}_{i}$. Let $\mathbf{Y}_{k}=\hat{M}(\mathbf{x}_{k})$, $\forall k\in[1:n]$. We first derive the conditional distribution of $\mathbf{Y}_{i}$ given $\{\mathbf{Y}_{j}$, $j\in\mathcal{U}_{all}\setminus\{\mathcal{U}_{col}\cup i\}\}$. Without loss of generality assume that $i=1$ and $\mathcal{U}_{col}=\{n-c+1,\dotsc,n\}$.
@@
The conditional distribution of $\mathbf{Y}_{1}$, given $\{\mathbf{Y}_{j}$, $j\in[2:n-c]\}$ is given by,
@@
$\tilde{\mathbf{\mu}}$ $\!=\!\mathbf{x}_{1}\!+\!\left(\!\!r\mathbf{1}_{n-c-1}^{T}\!\!\begin{pmatrix}\sigma^{2}+rc\!\!&\!\!r\!\!&\!\!\dotsc\!\!&\!\!r\\ r\!\!&\!\!\sigma^{2}+rc\!\!&\!\!\dotsc\!\!&\!\!r\\ \vdots\!\!&\!\!\vdots\!\!&\!\!\vdots\!\!&\!\!\vdots\\ r\!\!&\!\!r\!\!&\!\!\dotsc\!\!&\!\!\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\!\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)$
×$\qquad\qquad\qquad\times\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (78)
$\tilde{\Sigma}$ $\!=\!\!\left(\!\!\sigma^{2}\!+\!rc\!-\!r^{2}\mathbf{1}_{n\!-\!c\!-\!1}^{T}\!\!\begin{pmatrix}\sigma^{2}\!+\!rc\!\!&\!\!r\!\!&\!\!\dotsc\!\!&\!\!r\\ r\!\!&\!\!\sigma^{2}\!+\!rc\!\!&\!\!\dotsc\!\!&\!\!r\\ \vdots\!\!&\!\!\vdots\!\!&\!\!\vdots\!\!&\!\!\vdots\\ r\!\!&\!\!r\!\!&\!\!\dotsc\!\!&\!\!\sigma^{2}\!+\!rc\end{pmatrix}_{\!\!\!\!(n\!-\!c\!-\!1)^{2}}^{-1}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\mathbf{1}_{n\!-\!c\!-\!1}\!\!\right)\!\!\mathsf{I}_{d}$ (79)
@@
Proof:  [Proof of Claim 1] $\mathbf{Y}_{\ell}=\hat{M}(\mathbf{x}_{\ell})=\mathbf{x}_{\ell}+\sum_{j<\ell,j\notin\mathcal{U}_{col}}\mathbf{S}_{j,\ell}-\sum_{j>\ell,j\notin\mathcal{U}_{col}}\mathbf{S}_{\ell,j}+\mathbf{N}_{\ell}$, for all $\ell\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$. Therefore, $\mathbf{Y}_{\ell}\sim\mathcal{N}(\mathbf{x}_{\ell},(\sigma^{2}+rc)\mathsf{I}_{d})$. The distribution of the $k$th component of $\mathbf{Y}_{\ell}$, across all $\ell\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}=[1:n-c]$ is given by,
$\begin{pmatrix}Y_{1,k}\\ \vdots\\ \!Y_{n\!-\!c,k}\!\end{pmatrix}\!\!\sim\!\mathcal{N}\!\begin{pmatrix}\!\!\!\begin{pmatrix}x_{1,k}\\ \vdots\\ x_{n\!-\!c,k}\end{pmatrix}\!,\!\begin{pmatrix}\sigma^{2}\!+\!rc\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ r\!\!\!&\!\!\!\sigma^{2}\!+\!rc\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ \vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\\ r\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!\sigma^{2}\!+\!rc\!\!\!\!\!\end{pmatrix}_{\!\!\!(n\!-\!c)^{2}}\!\!\end{pmatrix}$ (81)
for $k\in[1:d]$, as $\text{cov}(Y_{i,k},Y_{j,k})=-\mathbb{E}[S_{i,j}^{2}]=r$ for $i\neq j$ and $\text{var}(Y_{i,k})=-r(n-c-1)+\sigma^{2}+r(n-1)=\sigma^{2}+rc$, $\forall i$. From the above distributions and from the fact that $\text{cov}(Y_{i,k},Y_{j,k^{\prime}})=0$, $i,j\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$ and $\forall k\neq k^{\prime}$, we derive,
$\begin{pmatrix}\mathbf{Y}_{1}\\ \vdots\\ \!\mathbf{Y}_{n\!-\!c}\!\end{pmatrix}\!\!\sim\!\mathcal{N}\!\begin{pmatrix}\!\!\!\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{n\!-\!c}\end{pmatrix}\!,\!\begin{pmatrix}(\sigma^{2}\!+\!rc)\mathsf{I}_{d}\!\!\!\!&\!\!\!\!r\mathsf{I}_{d}\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!r\mathsf{I}_{d}\\ r\mathsf{I}_{d}\!\!\!\!&\!\!\!\!(\sigma^{2}\!+\!rc)\mathsf{I}_{d}\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!r\mathsf{I}_{d}\\ \vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\\ r\mathsf{I}_{d}\!\!\!\!&\!\!\!\!r\mathsf{I}_{d}\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!(\sigma^{2}\!+\!rc)\mathsf{I}_{d}\!\!\!\!\!\end{pmatrix}\!\!\end{pmatrix}$ (82)
@@
Based on this, we have, $\mathbf{Y}_{1}|\mathbf{Y}_{2}\!=\!\mathbf{y}_{2},\dotsc,\mathbf{Y}_{n\!-\!c}\!=\!\mathbf{y}_{n\!-\!c}\sim N(\tilde{\mathbf{\mu}},\tilde{\Sigma})$ where,
$\tilde{\mathbf{\mu}}$ $=\mathbf{x}_{1}+(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})$
×$\times\left(\!\!\!\begin{pmatrix}(\sigma^{2}+rc)\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ r\!\!\!&\!\!\!(\sigma^{2}+rc)\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ \vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\\ r\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!(\sigma^{2}+rc)\end{pmatrix}\!\!\otimes\mathsf{I}_{d}\!\right)^{\!\!-1}\!\!\!\!\!\!\!\!\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n\!-\!c}\!-\!\mathbf{x}_{n\!-\!c}\end{pmatrix}$ (85)
$=\!\mathbf{x}_{1}\!+\!\!\!\left(\!r\mathbf{1}^{T}_{n\!-\!c\!-\!1}\begin{pmatrix}(\sigma^{2}+rc)\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ r\!\!\!&\!\!\!(\sigma^{2}+rc)\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ \vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\\ r\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!(\sigma^{2}+rc)\end{pmatrix}^{-1}\!\!\!\!\!\otimes\mathsf{I}_{d}\!\!\right)$
@@
$\tilde{\Sigma}$ $=(\sigma^{2}+rc)\mathsf{I}_{d}-(r\mathbf{1}_{n-c-1}^{T}\otimes\mathsf{I}_{d})$
×$\times\left(\!\!\!\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)^{2}}\!\!\!\!\!\!\!\!\!\!\!\otimes\mathsf{I}_{d}\right)^{-1}$